OpenStudy (anonymous):

The height of an equilateral triangular prism decreases by two units. The new volume is less than the original by how much? two times the area of the base two times the length of the height two more than the the area of the base two more than the length of the height

OpenStudy (eyust707):

Well, do you know the forumla for the area of a trianglular prism?

OpenStudy (anonymous):

v=Bh?

OpenStudy (inkyvoyd):

Yes.

OpenStudy (inkyvoyd):

What happens to the height? h-2 right?

OpenStudy (eyust707):

voume*

OpenStudy (anonymous):

B = area of base of the prism V prism = Bh So, say you start with Bh. Then, if the height shrinks by two... V new prism = B(h - 2) = Bh - 2B This is 2B less than the original, so the answer is two times the area of the base

OpenStudy (inkyvoyd):

Now put that in there, and compare it to the original answer.

OpenStudy (anonymous):

Thanks you so much guys!